QUESTION 16 Suppose 1 in 5 (1/5) JWU students own their own car. If four students are randomly selected, what is the probability that all four own their car? 。A. 1 /625 O B. 4/20 O C A/s O D. 1/20
Answer #1

Ans. A. 1/625

This is because if one in 5 owns their own car. We can say that each student has an independent probability of 1/5 to own their car

Probability that all four own their car = product.of probabilities of each one of them to own their car= ( 1/5)x (1/5)x (1/5) x (1/5) = 1/625.